OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3 ; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope = -3/5 ; y-intercept = (0, 2) slope = 5/3 ; y-intercept = (0, -6)

OpenStudy (anonymous):

rearrange the equation by solving for y, it will be: y = 1/3(5x+6) = (5/3)x +2 What is attached to the x is the slope and the other is the y intercept so 5/3x is slope and 2 is intercept First answer choice is correct

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome